1) In how many ways can a class of 30 students choose 4 of their representatives?
2) How many elements in the sample space of drawing a card out of a standard deck of 52, tossing a coin, and tossing a 12-face die?
3) You are drawing two cards out of a standard deck of 52 cards, without replacing the first one back, what is the probability of you drawing a face then an eight card?
I have a question, how do you know when to use Permutations or Combinations? Both seem similar so it's hard to figure out which equation to use :[
Thanks so much !

Answer by sudhanshu_kmr(1152)


1) no. of ways = 30C4 = 27405

2) number of sample space = 52 * 2 * 12 = 1248

3) probability = (12/52)* (4/51) = 4/221